OpenStudy (anonymous):

Which equation is solved by multiplying both sides by zero point five? z divided by zero point five equals thirty z plus zero point five equals thirty zero point five times z equals thirty z minus zero point five equals thirty

OpenStudy (anonymous):

z/0.5 = 30 multiply by 0.5 z = 45 z + 0.5 = 30 subtract by 0.5 z = 29.5 z * 0.5 = 30 divide by 0.5 z = 15